OpenStudy (anonymous):

Completely factor the following expression: 49y²– 81z²

OpenStudy (mayankdevnani):

(7y)^2-(9z)^2

OpenStudy (anonymous):

use \(a^2-b^2=(a+b)(a-b)\) with \(a=7y,b=9z\)

OpenStudy (mayankdevnani):

a^2-b^2=(a+b)(a-b)

OpenStudy (mayankdevnani):

(7y+9z)(7y-9z)

OpenStudy (mayankdevnani):

ok @sharron482

OpenStudy (anonymous):

thank you